Sprinkles285:

Alex and Lara have $21.00 each to spend at a book fair, where all students receive a 35% discount. They both want to purchase a copy of the same book, which normally sells for $28.50 plus 10% sales tax. To check if he has enough to purchase the book, Alex takes 35% of $28.50 and subtracts that amount from the normal price. He takes 10% of the discounted selling price and adds it back to find the purchase amount. Lara takes 65% of the normal purchase price and then computes 110% of the reduced price. Is Alex correct? Is Lara correct? Do they have enough money to purchase the book? Explain your answer using complete sentences and show your work. (10 points)

QueenAngel:

Hello :) Do you know how to solve this?

Sprinkles285:

nope

QueenAngel:

What would be the Total?

QueenAngel:

Okay so the purchase price of the book for students is: \[28.50 \times 1.10 \times 0.65 = ?\]

yarieliz:

Well you first The purchase price of the book for students is: 28.50*1.10*0.65 = $20.3775

yarieliz:

The you find Alex and Lara calculations Alex's calculation: (28.50*1.10 - 0.35*28.50) + 0.10*28.50*1.10*0.65 = Lara's calculation: 0.65*28.50*1.10 =

yarieliz:

From there you could get the answer and see which student is correct. I hope this helped.

Sprinkles285:

Thxs

yarieliz:

you welcome